T.H. Morgan also performed the following set of breeding experiments involving another mutant strain with yellow body colors, Xy, which is linked to X chromosome. The allele for the normal brown body color is indicated with Xy+. The interpretation of the breeding outcomes ended up discovering genetic linkage.

            P: female with normal eye and body color (w+,y+/w+,y+

                                    x

                 male with white eye and yellow body    (w,y/Y)

16. He also saw two minor categories, making up 1% of the progenies. Which were these two categories?

A. w ,y /Y and w+,y /Y

B. w+,y+/Y and w ,y /Y

C. w+,y /Y and w,y+/Y

D. w ,y+/Y and w+,y+/Y

E. w,y+/Y and w ,y /Y

17. These two minor categories are the products of __________ during the meiosis of F1 ______ flies.

A. segregation, male

B. segregation, female

C. crossing-over, male

D. crossing-over, female

E. independent assortment, female

18. Why didn't he count the number of F2 female flies?

A. all the F2 female flies were normal eye color and body color.

B. all the F2 female flies showed the exact same patterns as the male flies.

C. all the F2 females showed unexplainable outcomes.

D. all of the above.

E. none of the above

ONLY NEED #18 ANSWERED THE # 16 AND 17 ARE FOR CONTEXT
Answer #1

18. Option A is correct
Explanation:
Parental cross: w+y+/w+y+ X w,y/Y
F1 progeny: w+y+/wy w+y+/Y
F1 selfing: w+y+/wy w+y+/Y
F2 progeny:
w+y+/w+y+ = Female = Normal eye color and body color
wy/w+y+ = Female = Normal eye color and body color
w+y+/Y = Male = Normal eye color and body color
wy/Y = Male = Mutant eye color and body color

The above phenotypic classes will be optained assuming complete linkage.i.e. no recombination between w+ and y+ genes.